A jet plane lands with a speed of 100m/s when it touches the runway and can accelerate uniformly at a maximum rate of -5.0 m/s² as it comes to rest. Can this plane land at an airport where the runway is 800 m long?
1
Expert's answer
2020-09-16T10:09:40-0400
"v_f^2-v_i^2=2as\\v_f^2-100^2=2(-5)(800)\\v_f=44.7\frac{m}{s}>0\frac{m}{s}"

No. This is too short stopping distance.